OpenStudy (anonymous):

Suppose f and g are continuous functions such that g(8) = 2 and the limit as x approaches 8 [3f(x) + f(x)g(x)]= 30. Find f(8). Do you treat f(8) as if it is a variable and solve for it?

OpenStudy (anonymous):

yes. you have 3f(8) + f(8)g(8) = 30 3f(8) + 2f(8) = 30 5f(8) = 30

OpenStudy (anonymous):

okay so i got that but now im confused i was given this question: Suppose f and g are continuous functions such that g(4) = 2 and lim x → 4 [3f(x) + f(x)g(x)] = 25. Find f(4). Im stuck at this step: 25=[3f(4) + f(4)(2)] Help Me!!

OpenStudy (anonymous):

3f(4) + 2f(4) = 5f(4)

OpenStudy (anonymous):

I never understood that step. seeing it really helped. your awesome, Thank You. ^_^

OpenStudy (anonymous):

glad i could help :)